The middle school team score 3 field goals worth three points each and two touchdowns with extra points worth seven points each. Write a numerical expression to find the team's score. Then evaluate the expression. Show your work.
Answer by stanbon(75887)

The middle school team score 3 field goals worth three points each and two touchdowns with extra points worth seven points each. Write a numerical expression to find the team's score. Then evaluate the expression. Show your work.
Points = 3f + 7t
------------------
If f = 3 and t = 2, points = 3*3 + 2*7 = 9+14 = 23
----------------
Cheers,
Stan H.
---------------